# Micellization in active matter of asymmetric self-propelled particles: Experiments

### v2

@@ IV.3 Micellization in swarms of non-chiral robots in a parabolic potential
We proceed with studying a system of non-chiral -robots moving with low activity ()[Fig. 5(a)] and placed in a parabolic satellite dish having the dimensions of [Fig. 1(a)]. In contrast to a hard wall, such a dish creates a soft localizing potential and, as we will demonstrate further, prevents the condensation of robots at the boundary. The micellization parameter fluctuates during the system evolution but does not demonstrate any pronounced growth characteristic of micellization, Fig. 5(a). In this case, robots are predominantly directed away from the vertex of the parabolic potential Dauchot and Démery (2019). Covering the lateral surfaces of robots with abrasive tape in order to increase the friction between them does not result in any systematic changes of the micellization parameter, see the blue solid line in Fig. 5(a).
Next, we consider the dependencies of the micellization parameter on the activity of the robots, their center of inertia location, and the packing density that is varied by changing the number of robots in the parabolic potential. For systems of -robots with low activity () shown in Fig. 6(a), the micellization parameter fluctuates near the minimal level () independently of the density of the system, and micelle formation is not observed; see Table 3. Increasing the activity to results in the instability of the robot’s motion at higher densities . In particular, the robots begin to flip, changing the number of self-propelled particles in the system. The corresponding curves of the micellization parameter are not shown; see the gray regions in Fig. 6(b),(f).
In contrast to -robots, -robots do not flip even at high activity. They do not demonstrate any micelles as well in the experiments with sparse swarms of robots. However, stable half-micelles appear in these systems, characterized by an increase in the micellization parameter (). The formation of transient micelles with a short lifetime of a few seconds emerges at the higher density , and is well expressed for , indicated by multiple peaks of the micellization parameter in Fig. 6(c),(d) both for low () and high () activities. Increasing robot activity results in shorter average micelle lifetimes, namely s for and s for for high robot activity, compared to s for and s for for low activity.
Next, we modify the system by covering the lateral surfaces of all robots with an abrasive (P800 sandpaper) Sup to consider how the micellization changes in the case of high friction between the particles, which may prove important for microscale implementations. As seen in Fig. 6(e),(f), the behavior of -robots is similar to the case of particles with low lateral friction, Fig. 6(a),(b). The micellization parameter slightly fluctuates, indicating the absence of micellization, and systems with high activity () and become unstable. The same is confirmed by the data in Table 3.
@@
However, for -robots with abrasive lateral surfaces, the situation becomes strikingly different compared to the case with smooth lateral surfaces. As seen in Table 3, micelle formation for low activity () is observed only for the system with , while for sparse swarms with or, in contrast, dense swarms with the micelles are not observed. Moreover, for high activity (), micelles are observed most frequently for intermediate density , while for micelle formation is less pronounced and appears almost completely suppressed for , with only a single experiment featuring the formation of a complete micelle. Instead, distorted micelle-like clusters consisting of two half-micelles shifted a few centimeters relative to each other frequently form in these systems, characterized by plateaus of the micellization parameter with values ; see Fig. 6(g),(h). The average lifetimes of micelles are s for (excluding the anomalously stable micelle discussed separately in Sec. VI), s for and s for and significantly exceed those of the micelles formed by smooth robots. Decreasing robot activity leads to an even longer average micelle lifetime of s for . Thus, it is seen that, in contrast to robots with smooth lateral surfaces, micelles are more stable in swarms of abrasive robots, and increasing robot activity significantly enhances the probability of micelle formation.
@@
This difference in micellization between smooth and abrasive robots could be related to the improved stability of the clusters formed by robots with high lateral friction. Systems of robots with smooth lateral surfaces demonstrate fewer micelles at lower densities, while at higher densities there are more micelles that decompose quickly. This might be explained by the increased probability of micelle formation during the decomposition of other clusters, which are generally characterized by low stability. With high lateral friction, micelles demonstrate longer lifetimes, apparently related to restriction of the relative lateral movement of adjacent robots in a cluster. This would presumably lead to increased micellization in systems forming clusters that are more likely to merge into micelles, which, once formed, would remain stable. In contrast, for systems where disordered clusters are more likely to form, their enhanced stability would lower the chances of decomposition followed by micelle formation, resulting in decreased micellization. Such a variety of possible mechanisms emphasizes an complex interplay between the asymmetry of individual particles, their activity, and friction in the considered system. Moreover, the fact that micellization is more prominent for higher lateral friction points towards the possibility of microscale realizations, for example, using Janus particles Zhang et al. (2015); Hils et al. (2021).
V Experimental studies of microscopic micellization mechanisms
@@ V.1 Similar single-particle dynamics of robots with different center of inertia locations
We start by considering the motion of a single self-propelled robot placed in a parabolic potential implemented with the help of a satellite dish, the same as in the swarm experiments discussed in Section IV.3. The dish has two different main axes cm and cm, and therefore its horizontal cross sections are ellipses. Such a soft localizing potential allows us to study the robot’s motion for an arbitrary duration, with the robots neither escaping the area captured by the camera nor condensing at the boundary.
@@
In our experiments, we vary the robot activity (defined by PWM), the center of inertia by setting it at points and , and the initial robot placement configuration defined by the robot’s position and angular orientation. PWM takes the values (the low activity) and (the high activity) corresponding to the robot motion velocities of approximately cm/s and cm/s for -robots, and cm/s and cm/s for -robots, respectively. Four initial robot placement configurations are considered: (i) the “center” configuration corresponding to the robot placed at the vertex of the parabolic potential and its propulsion vector oriented along the semi-minor axis, Fig. 7(a); (ii) the “edge” configuration with the robot placed at the distance cm from the vertex at the semi-minor axis and the propulsion vector directed towards the vertex point, Fig. 7(b); (iii) the “orbit CCW” configuration with the robot placed at the semi-minor axis at the distance cm from the vertex and the propulsion vector directed along the semi-major axis in such a manner that the robot moves in a CCW-direction, Fig. 7(c); and (iv) the “orbit CW” configuration when the robot is placed at the distance cm from the vertex at the semi-minor axis and the propulsion vector is directed along the semi-major axis resulting in a CW-directed robot’s motion, Fig. 7(d). For each set of parameters, we repeated the experiment with three different robots whose motion has been recorded during seconds. To study the kinematic characteristics, we introduce the radius vector of robot , defined in polar coordinates and with the origin at the vertex of the parabolic potential and the angle measured from the major semi-axis of the parabolic satellite dish. The radius vector points to the center of the robot’s marker. We also introduce the robot orientation angle between the robot propulsion vector and the radius vector.
@@
Figure 7(e)-(l) demonstrates the time evolution of the radius vector components and the orientation angle of a single robot in the parabolic potential. The four previously described initial placement configurations of the robot moving with high activity () are studied for an -robot and an -robot with results displayed in Fig. 7(e)-(h) and Fig. 7(i)-(l), respectively. It is seen that there are two major types of individual robot motion, in accordance with Dauchot and Démery (2019): a car drifting-like climbing motion whose distinctive feature is the propulsion vector directed outwards the vertex of the parabolic potential and the average velocity directed tangentially to height isolines, and the orbiting motion with the propulsion vector oriented tangentially to the elliptical trajectory of the robot.
@@
The first type of motion is characterized by the orientation angle close to , see the dashed horizontal lines in Fig. 7(e)-(l). Moreover, the radial coordinate and the polar angle change slowly (up to some fast fluctuations), with most of the curves corresponding to the robot completing a single revolution around the potential’s vertex during the observed timespan [see the blue, green, and orange solid lines in Fig. 7(e) and the green solid line in Fig. 7(g),(k),(l)] or even an incomplete rotation [the orange solid lines in Fig. 7(h) and Fig. 7(i)-(l) and the green solid lines in Fig. 7(i),(j)]. The radial coordinate changes with the same oscillation period as corresponding to the robot moving at the same height along the elliptical trajectory characterized by different values of the semi-minor and semi-major axes. Due to the ellipticity of the trajectory, oscillations with the same period can also be observed in the orientation angle , which would not exist if the trajectory was a perfect circle. The described climbing motion occurs for the -robot, Fig. 7(e)-(h) [the blue solid line in Panel (e), the green solid lines in Panels (e) and (g), and the orange solid lines in Panels (e) and (h)] as well as for the -robot [the green and orange solid lines in Fig. 7(i)-(l)]. The fast fluctuations mentioned previously, which are clearly manifested in the radial coordinate and the orientation angle , are related to the robot’s oscillating up-and-down motion around a steady height level, see Supplementary Video 4. In addition, the rattling of with frequencies greatly exceeding those of the oscillations of and an amplitude that may in some cases reach [see Fig. 7(e)] results from a noise in the robot’s propulsion vector orientation due to the bristle-bot vibrating motion mechanism. It is seen that such rattling is only present in the orientation angle but absent in the polar angle .
@@
The second type of single robot motion is characterized by fast oscillations of the radial coordinate accompanied by a sawtooth waveform in the polar angle , both having the same period, Fig. 7. This signifies that the robot repeatedly traverses an elliptical trajectory, while the period of these oscillations is inversely proportional to the robot’s tangential velocity. Moreover, oscillations with a more complicated waveform are observed in the orientation angle that stays either in the vicinity of or , indicating that the robot’s propulsion vector direction is well aligned with its direction of motion along the elliptical trajectory. Thus, in contrast to the climbing motion described above, this second type of motion resembles the movement of a car running on an oval track by proper steering. Similarly to the previous motion type, this orbiting motion is observed for both the -robots shown with the blue and orange solid lines in Fig. 7(f),(g) and the blue solid line in Fig. 7(h), and the -robots shown with the blue solid lines in Fig. 7(i)-(l).
@@
In some cases, a hybrid type of a single robot motion can be observed, characterized by the orientation angle taking the values between and and the robot making between two and three complete revolutions around the vertex per minute; see the solid green lines in Fig. 7(f)-(h). The presence of such motion profiles with intermediate characteristics demonstrates that the two previously described regimes may be considered as limiting cases of a single type of motion corresponding to the robot moving periodically along an elliptical trajectory with a certain angle between the robot’s velocity vector and its propulsion vector. Greater values of this angle correspond to more pronounced radial oscillations with a longer period.
@@ V.2 Dominant stable two-particle cluster type defined by particle center of inertia location
To address microscopic micellization mechanisms beyond the single-particle level, we proceed with considering the collisions of two self-propelled robots for different sets of parameters: low () and high () activity of the robots, abrasive and smooth lateral surfaces, and impact angles ; see Fig. 8(a). Despite single-robot experiments were conducted in a shallow parabolic potential of a satellite dish, we consider non-chiral robots moving linearly on a plane when studying two-robot collisions, as controlling the impact angle for robots moving along elliptical or circular trajectories appears extremely challenging.
@@
Figure 8 shows the key results for - and -robots with smooth and abrasive lateral surfaces moving with high activity and colliding at impact angles [Fig. 8(a)-(d)] and [Fig. 8(e)-(h)] while additional data for low activity, as well as for are considered in the Supplemental Material Sup . For each set of parameters, we perform three to four experiments with different pairs of robots to ensure that the observed properties of scattering and cluster formation indeed represent the characteristic behavior of the considered system rather than the specific properties of certain robots. In the following, we focus on the evolution of the distance between the noses of the robots and the angle between their propulsion vectors.
@@
It is clearly observed that, regardless of the impact angle and the presence of an abrasive coating on their lateral surfaces, -robots form stable (i.e. moving without any considerable shape variations) heart-shaped two-robot clusters upon a collision, see the inset in Fig. 8(b) and Supplementary Video 5. As seen in Fig. 8(b),(d),(f),(h), formation of such clusters is indicated by an almost vanishing distance between the robots’ noses and the angle between the robots’ propulsion vectors tending to (the gray dashed lines in Fig. 8). In particular, during the first stage (lasting approximately s) of the experiment, the angle keeps its initial value , while the distance between robots’ noses decreases linearly. After that, the second stage of a collision takes place, hallmarked by an abrupt change in the angle from to accompanied by the distance reaching cm. Physically, this corresponds to the robots adjusting their propulsion vector directions by rotating around the point of contact until is reached. Finally, during the third stage, both and cm remain nearly unchanged (up to some noise-like fluctuations), highlighting that two robots now propel forward in a continuous contact. The described “heart” clusters essentially represent quarter-micelles and can potentially contribute to the formation of half-micelles (consisting of four robots) upon a collision between two such clusters.
@@
The described scenario of “heart” cluster formation is most clearly observed in Fig. 8(b),(d) for the robots starting with which considerably differs from , while for the transition region is much less pronounced, as seen in Fig. 8(f),(h). However, due to the increased technical difficulty in engineering the collisions between the robots for compared to , there are considerable deviations between the target value of and the actual impact angle obtained. In particular, Fig. 8(h) demonstrates close to in all three experiments. Moreover, in some cases collisions of the robots do not result in cluster formation [see the blue and orange solid lines in Fig. 8(h)].
@@
In contrast, for -robots the “heart” clusters are unstable, as seen in Fig. 8(a),(c). In particular, when and reach their values of and cm characteristic of the heart-shaped clusters, they fluctuate around their mean values for s. After that, starts growing until it reaches cm, while decreases to , as seen for all experiments in Fig. 8(a) and for the experiment shown with the green solid line in Fig. 8(c). The meaning of these characteristic values of and is uncovered in Fig. 8(e). It turns out that -robots support the formation of a different stable cluster type – the “tandem” cluster – in which one robot touches with its nose the round part of the other robot, as shown in the inset in Fig. 8(e) and Supplementary Video 5. Such clusters can also travel with the values of and retained, up to some fluctuations. Thus, the “heart” cluster in the case of -robots is inherently metastable and eventually transforms into the asymmetric “tandem” cluster described above. Note that for higher values of the impact angle , the formation of such “tandem” clusters is more pronounced for robots with abrasive lateral surfaces, as seen from the comparison of Fig. 8(e) and Fig. 8(g).

### v3

@@ IV.3 Micellelike clusters in swarms of nonchiral robots in a parabolic potential
However, for -robots with abrasive lateral surfaces, the situation becomes strikingly different compared to the case with smooth lateral surfaces. As seen in Table 3, micelle formation for low activity () is observed only for the system with , while for sparse swarms with or, in contrast, dense swarms with the micelles are not observed. Moreover, for high activity (), micelles are observed most frequently for intermediate density , while for micelle formation is less pronounced and appears almost completely suppressed for , with only a single experiment featuring the formation of a complete micelle. Instead, distorted micelle-like clusters consisting of two half-micelles shifted a few centimeters relative to each other frequently form in these systems, characterized by plateaus of the micellization parameter with values ; see Figs. 6(g) and 6(h). The average lifetimes of micelles are s for (excluding the anomalously stable micelle discussed separately in Sec. VI), s for , and s for , significantly exceeding those for the micelles formed by smooth robots. Decreasing robot activity leads to an even longer average micelle lifetime of s for . Thus, it is seen that, in contrast to robots with smooth lateral surfaces, micelles are more stable in swarms of abrasive robots, and increasing robot activity significantly enhances the probability of micelle formation.
@@
This difference in micellization between smooth and abrasive robots could be related to the improved stability of the clusters formed by robots with high lateral friction. Systems of robots with smooth lateral surfaces demonstrate fewer micelles at lower densities, while at higher densities there are more micelles that decompose quickly. This might be explained by the increased probability of micelle formation during the decomposition of other clusters, which are generally characterized by low stability. With high lateral friction, micelles demonstrate longer lifetimes, apparently related to restriction of the relative lateral movement of adjacent robots in a cluster. This would presumably lead to increased micellization in systems forming clusters that are more likely to merge into micelles, which, once formed, would remain stable. In contrast, for systems where disordered clusters are more likely to form, their enhanced stability would lower the chances of decomposition followed by micelle formation, resulting in decreased micellization. Such a variety of possible mechanisms emphasizes the complex interplay between the asymmetry of individual particles, their activity, and friction in the considered system, which may be crucial for microscale realizations, for example, using Janus particles Zhang et al. (2015); Hils et al. (2021).
To evaluate the interaction of micellelike clusters with nonmicellar aggregates, lifetimes of manually assembled micellelike clusters were assessed. First, manually assembled micellelike clusters in the absence of nonmicellized robots were examined. They remained stable for the entire duration of the experiment (several minutes), independently of the vibration activity of the robots (low or high), their lateral friction (smooth or abrasive side surfaces), and even of the center of inertia location-micellelike clusters composed of both -robots and -robots remained stable (see supplementary video 7). From this we can conclude that micellization is governed by the probability of micellelike cluster formation, but once formed, micellelike clusters are stabilized by the parabolic potential. After that, nonmicellized robots were added to evaluate the interaction of micellelike clusters with other aggregates and individual robots. These systems most often displayed two kinds of behavior: micellelike clusters either decomposed in the first few seconds after a collision with surrounding robots, or remained stable for the entire experiment if such collisions did not occur (see supplementary video 8). Additionally, -robots occasionally formed new micellelike clusters, while -robots did not (see the supplementary material Sup for a detailed analysis). Therefore, micellelike clusters continuously form and decompose with probabilities governed by the interactions between individual robots and various clusters.
V Microscopic mechanisms of micelle-like clusters formation
@@ V.1 Similar single-particle dynamics of robots with different center of inertia locations
We start by considering the motion of a single self-propelled robot placed in a parabolic potential implemented with the help of a satellite dish, the same as in the swarm experiments discussed in Sec. IV.3. The dish has two different main axes cm and cm, and therefore its horizontal cross sections are ellipses. Such a soft localizing potential allows us to study the robot’s motion for an arbitrary duration, with the robots neither escaping the area captured by the camera nor condensing at the boundary.
@@
In our experiments, we vary the robot activity (defined by PWM), the center of inertia by setting it at points and , and the initial robot placement configuration defined by the robot’s position and angular orientation. PWM takes the values (the low activity) and (the high activity) corresponding to the robot motion velocities of approximately and for -robots, and cm/s and cm/s for -robots, respectively. Four initial robot placement configurations are considered: (i) the “center” configuration corresponding to the robot placed at the vertex of the parabolic potential and its propulsion vector oriented along the semiminor axis, Fig. 7(a); (ii) the “edge” configuration with the robot placed at the distance cm from the vertex at the semiminor axis and the propulsion vector directed towards the vertex point, Fig. 7(b); (iii) the “orbit CCW” configuration with the robot placed at the semi-minor axis at the distance cm from the vertex and the propulsion vector directed along the semimajor axis in such a manner that the robot moves in a CCW-direction, Fig. 7(c); and (iv) the “orbit CW” configuration when the robot is placed at the distance cm from the vertex at the semi-minor axis and the propulsion vector is directed along the semimajor axis resulting in a CW-directed robot’s motion, Fig. 7(d). For each set of parameters, we repeated the experiment with three different robots whose motion has been recorded during s. To study the kinematic characteristics, we introduce the radius vector of robot , defined in polar coordinates and with the origin at the vertex of the parabolic potential and the angle measured from the major semiaxis of the parabolic satellite dish. The radius vector points to the center of the robot’s marker. We also introduce the robot orientation angle between the robot propulsion vector and the radius vector.
@@
Figures 7(e)-7(l) demonstrates the time evolution of the radius vector components and the orientation angle of a single robot in the parabolic potential. The four previously described initial placement configurations of the robot moving with high activity () are studied for an -robot and an -robot with results displayed in Figs. 7(e)-7(h) and Figs. 7(i)-7(l), respectively. It is seen that there are two major types of individual robot motion, in accordance with Ref. Dauchot and Démery (2019): a car drifting-like climbing motion, whose distinctive feature is the propulsion vector directed outwards the vertex of the parabolic potential and the average velocity directed tangentially to height isolines, and the orbiting motion with the propulsion vector oriented tangentially to the elliptical trajectory of the robot.
@@
The first type of motion is characterized by the orientation angle close to ; see the dashed horizontal lines in Figs. 7(e)-7(l). Moreover, the radial coordinate and the polar angle change slowly (up to some fast fluctuations), with most of the curves corresponding to the robot completing a single revolution around the potential’s vertex during the observed timespan [see the blue, green, and orange solid lines in Fig. 7(e) and the green solid line in Figs. 7(g), 7(k), and 7(l)] or even an incomplete rotation [the orange solid lines in Fig. 7(h) and Figs. 7(i)-7(l) and the green solid lines in Figs. 7(i) and 7(j)]. The radial coordinate changes with the same oscillation period as corresponding to the robot moving at the same height along the elliptical trajectory characterized by different values of the semiminor and semimajor axes. Due to the ellipticity of the trajectory, oscillations with the same period can also be observed in the orientation angle , which would not exist if the trajectory was a perfect circle. The described climbing motion occurs for the -robot, Figs. 7(e)-7(h) [the blue solid line in Fig. 7(e), the green solid lines in Figs. 7(e) and 7(g), and the orange solid lines in Figs. 7(e) and 7(h)] as well as for the -robot [the green and orange solid lines in Figs. 7(i)- 7(l)]. The fast fluctuations mentioned previously, which are clearly manifested in the radial coordinate and the orientation angle , are related to the robot’s oscillating up-and-down motion around a steady height level, see supplementary video 4. In addition, the rattling of with frequencies greatly exceeding those of the oscillations of and an amplitude that may in some cases reach [see Fig. 7(e)] results from a noise in the robot’s propulsion vector orientation due to the bristle-bot vibrating motion mechanism. It is seen that such rattling is only present in the orientation angle but absent in the polar angle .
@@
The second type of single robot motion is characterized by fast oscillations of the radial coordinate accompanied by a sawtooth waveform in the polar angle , both having the same period, Fig. 7. This signifies that the robot repeatedly traverses an elliptical trajectory, while the period of these oscillations is inversely proportional to the robot’s tangential velocity. Moreover, oscillations with a more complicated waveform are observed in the orientation angle that stays either in the vicinity of or , indicating that the robot’s propulsion vector direction is well aligned with its direction of motion along the elliptical trajectory. Thus, in contrast to the climbing motion described above, this second type of motion resembles the movement of a car running on an oval track by proper steering. Similarly to the previous motion type, this orbiting motion is observed for both the -robots shown with the blue and orange solid lines in Figs. 7(f) and 7(g) and the blue solid line in Fig. 7(h), and the -robots shown with the blue solid lines in Figs. 7(i)-7(l).
@@
In some cases, a hybrid type of a single robot motion can be observed, characterized by the orientation angle taking the values between and and the robot making between two and three complete revolutions around the vertex per minute; see the solid green lines in Figs. 7(f)-7(h). The presence of such motion profiles with intermediate characteristics demonstrates that the two previously described regimes may be considered as limiting cases of a single type of motion corresponding to the robot moving periodically along an elliptical trajectory with a certain angle between the robot’s velocity vector and its propulsion vector. Greater values of this angle correspond to more pronounced radial oscillations with a longer period.
To evaluate the role of collisions in the single-particle motion of robots, rotational excitation was applied in the form of instantaneous strikes to various points of the robot side surface far from its center of inertia. The strike causes a transient rotation, which damps in s for -robots and s for -robots. After that, the robots return to their normal movement pattern, that is, climbing or orbiting motion. Detailed analysis of these experiments is provided in the supplementary material Sup . Such damping times are significantly smaller than other characteristic times in the system. Therefore, transport of angular momentum can be neglected due to its almost instantaneous damping by bristle-bots, and robot collisions that do not result in cluster formation are unlikely to play a significant role in micellization.
@@ V.2 Dominant stable two-particle cluster type defined by particle center of inertia location
To address microscopic micellization mechanisms beyond the single-particle level, we proceed with considering the collisions of two self-propelled robots for different sets of parameters: low () and high () activity of the robots, abrasive and smooth lateral surfaces, and impact angles ; see Fig. 8(a). Despite single-robot experiments were conducted in a shallow parabolic potential of a satellite dish, we consider nonchiral robots moving linearly on a plane when studying two-robot collisions, as controlling the impact angle for robots moving along elliptical or circular trajectories appears extremely challenging.
@@
Figure 8 shows the key results for - and -robots with smooth and abrasive lateral surfaces moving with low and high activities and colliding at impact angles [Figs. 8(a)-8(d)] and [Figs. 8(e)-8(h)] while additional data for low activity, as well as for are considered in the supplementary material Sup . For each set of parameters, we perform three to four experiments with different pairs of robots to ensure that the observed properties of scattering and cluster formation indeed represent the characteristic behavior of the considered system rather than the specific properties of certain robots. In the following, we focus on the evolution of the distance between the noses of the robots and the angle between their propulsion vectors.
@@
It is clearly observed that, regardless of the impact angle and the presence of an abrasive coating on their lateral surfaces, -robots form stable (i.e., moving without any considerable shape variations) heart-shaped two-robot clusters after a collision,; see the inset in Fig. 8(b) and supplementary video 5. As seen in Figs. 8(b), 8(d), 8(f),and 8(h), formation of such clusters is indicated by an almost vanishing distance between the robots’ noses and the angle between the robots’ propulsion vectors tending to (the gray dashed lines in Fig. 8). In particular, during the first stage (lasting approximately s) of the experiment, the angle keeps its initial value , while the distance between robots’ noses decreases linearly. After that, the second stage of a collision takes place, hallmarked by an abrupt change in the angle from to accompanied by the distance reaching cm. Physically, this corresponds to the robots adjusting their propulsion vector directions by rotating around the point of contact until is reached. Finally, during the third stage, both and cm remain nearly unchanged (up to some noise-like fluctuations), highlighting that two robots now propel forward in a continuous contact. The described “heart” clusters essentially represent quarter-micelles and can potentially contribute to the formation of half-micelles (consisting of four robots) after a collision between two such clusters.
@@
The described scenario of “heart” cluster formation is most clearly observed in Figs. 8(b) and 8(d) for the robots starting with which considerably differs from , while for the transition region is much less pronounced, as seen in Figs. 8(f) and 8(h). However, due to the increased technical difficulty in engineering the collisions between the robots for compared to , there are considerable deviations between the target value of and the actual impact angle obtained. In particular, Fig. 8(h) demonstrates close to in all three experiments. Moreover, in some cases collisions of the robots do not result in cluster formation [see the blue and orange solid lines in Fig. 8(h)].
@@
In contrast, for -robots the “heart” clusters are unstable, as seen in Figs. 8(a) and 8(c). In particular, when and reach their values of and cm characteristic of the heart-shaped clusters, they fluctuate around their mean values for s. After that, starts growing until it reaches cm, while decreases to , as seen for all experiments in Fig. 8(a) and for the experiment shown with the green solid line in Fig. 8(c). The meaning of these characteristic values of and is uncovered in Fig. 8(g). It turns out that -robots support the formation of a different stable cluster type – the “tandem” cluster – in which one robot touches with its nose the round part of the other robot, as shown in the inset in Fig. 8(g) and supplementary video 5. Such clusters can also travel with the values of and retained, up to some fluctuations. Thus, the “heart” cluster in the case of -robots is inherently metastable and eventually transforms into the asymmetric “tandem” cluster described above. Note that for higher values of the impact angle , the formation of such “tandem” clusters is more pronounced for robots with abrasive lateral surfaces, as seen from the comparison of Fig. 8(e) and Fig. 8(g).